# A time warping model for seasonal data with application to age estimation from narwhal tusks

### v1

@@ Abstract
Signals with varying periodicity frequently appear in real-world phenomena, necessitating the development of efficient modelling techniques to map the measured nonlinear timeline to linear time. Here we propose a regression model that allows for a representation of periodic and dynamic patterns observed in time series data. The model incorporates a hidden strictly increasing stochastic process that represents the instantaneous frequency, allowing the model to adapt and accurately capture varying time scales. A case study focusing on age estimation of narwhal tusks is presented, where cyclic element signals associated with annual growth layer groups are analyzed. We apply the methodology to data from one such tusk collected in West Greenland and use the fitted model to estimate the age of the narwhal. The proposed method is validated using simulated signals with known cycle counts and practical considerations and modelling challenges are discussed in detail. This research contributes to the field of time series analysis, providing a tool and valuable insights for understanding and modeling complex cyclic patterns in diverse domains.
@@
stochastic growth-time process,
@@ 1 Introduction
In some applications, the measurements are taken along a segment, where each measured point at some distance corresponds to an unknown time. Key examples are tree rings, sea sediment cores, ice cores, hair, ear stones (Monserud and Marshall, 2001; Jouzel and Masson-Delmotte, 2010; Charles et al., 1996; Hüssy et al., 2021a; Mosbacher et al., 2016), or in this case, narwhal tusks. A given variable is measured along a transect line at equidistant spatial points, and the oscillation patterns in the data are hypothesized to reflect yearly variations. The measurements can then be dated, i.e., distances can be transformed to a time line.
@@
A significant challenge presented by such datasets relates to the unknown nonlinear relation between distance and the underlying timeline, complicating the precise mapping of the signals to their actual temporal positions. In cases where reference signals are available or when the primary objective of the study centers on signal alignment rather than precise temporal mapping, methodologies like Graphical and Dynamic Time Warping (Jeong, Jeong and Omitaomu, 2011; Salvador and Chan, 2007; Wang et al., 2016) offer effective solutions for tasks involving pattern matching and classification. However, when reference signals are absent, and the exact mapping from spatial position to real time is of particular concern, it appears that no universal method for time warping exists, to the best of our knowledge.
@@
This paper presents a novel approach for modelling noisy non-stationary cyclic signals with changing periodicity and amplitude. As opposed to the traditional seasonal decomposition (Dagum, 2013; Cleveland et al., 1990) where a signal is split into trend, seasonal, cyclic, and noise components, our model incorporates the seasonal and cyclic effects, and combine them into a single component. A key innovation in our model is the incorporation of a hidden strictly positive stochastic process that models the (instantaneous) growth process of the material. The integral of this process is strictly increasing and models the time-growth process, which enables the modulation of frequencies while respecting the direction of time. This provides realism in capturing the varying periodicity observed in real-world data. Section 2 outlines the model, while Section 3 explains the estimation process. In section 4, we simulate from our model.
@@
The flexibility of our model allows for a wide range of applications. For example, in ecological studies, time series data are often associated with annual cycles, such as the growth patterns of organisms where various exogenous and endogenous factors can influence the speed of growth (Stounberg et al., 2022; Amais et al., 2021; Wooller et al., 2021). Our method can effectively capture the changing frequencies and provide more accurate predictions for these cyclic patterns.
Narwhals (Monodon monoceros), a toothed whale endemic to the Arctic regions, exhibit annual depositions of growth layer groups (GLGs) in their tusks (Garde et al., 2024) resulting in annual cyclic patterns. We present a case study in section 5 where we demonstrate and validate our model on element signals (specifically the isotope Barium-137) derived from one narwhal tusk, and use the method to estimate the number of cycles, thereby also the age of the narwhal. We conclude with a discussion of the model in Section 6.
@@ 2 Time warping model
Consider a time series with an additive decomposition,
$y_{i}=f(x_{i},\theta)+\epsilon_{i},\quad i=0,1,\ldots,n;\,x_{i}=i\Delta$ (1)
where $y_{i}$ is the measured variable (e.g., a chemical element) at spatial location $x_{i}$ and $\Delta=x_{i+1}-x_{i}$ is the distance between measurements. The observation interval is thus $[0,n\Delta]$. The function $f$ is detailed below and depends on parameters $\theta$. The error terms $\epsilon_{i}$ are assumed independent and normally distributed with mean 0 and variance $\sigma^{2}$. The variable $x$ is the distance along a cross section of some object of interest, where a specific position corresponds to an unknown time through a monotonic and possibly non-linear function $g$,
$t_{i}=g(x_{i},\theta).$ (2)
We impose that $g(\cdot)$ is a strictly increasing function with respect to $x$, to ensure that time is not going backwards.
For the signal, we assume the form
$f(x,\theta)=A\sin(g(x,\theta)+b)-B(x)\cos(2g(x,\theta)+2b),\quad g(0,\theta)=0,$ (3)
where $B(x)>A>0$ determine the amplitudes, which are allowed to vary over time through $B(x)$, and $b$ is the phase at $x=0$. The condition $B(x)>A$ ensures that the lower and upper envelopes of the signal do not intersect. In eq. (3) there is no trend, which we assume is removed before the analysis, if needed. Model (3) is a truncated harmonic regression model using only two harmonics, and where $(B(x)+A)$ and $(B(x)-A)$ are local maxima, which we denote as the high and low peaks, respectively. Thus, the observations are composed of a slow wave, a fast wave (twice the speed of the slow wave) and measurement noise. The slow wave can for example be an annual cycle, and the fast wave is then summer-winter differences.
In the linear and deterministic case where $g(x,\theta)=ax$ the cycles are regularly occurring and the decomposition is known as a seasonal decomposition (Cleveland et al., 1990). However, this is unrealistic in most settings, where the length of one cycle will typically vary due to environmental factors. We therefore allow the instantaneous growth rate to be stochastic. We denote the unknown stochastic transformation $g(x,\theta)$ by the growth-time process.
[…]
@@
where $W_{x}$ is a standard Wiener process. The hidden process $(\xi_{x})_{x\geq 0}$ describes instantaneous frequency deviations from the baseline frequency $a/2\pi$. The parameter $\beta$ determines the rate of adjustment to the baseline and $\omega$ scales the noise level. This is an ergodic process with stationary distribution the Gamma distribution with shape parameter $2\beta a/\omega^{2}$, scale parameter $\omega^{2}/(2\beta)$, mean $a$ and variance $a\omega^{2}/2\beta$, provided that $\beta,a,\omega>0$, and $2\beta a\geq\omega^{2}$ (Ditlevsen, Rubio and Lansky, 2020). These parameter restrictions ensure that any solution to Eq. (4) is strictly positive (except at $x=0$) with long-term mean $a$, leading to a strictly increasing process $g(x,\theta)$. Define also the one-lag autocorrelation of the growth process $\rho=\exp(-\Delta\beta)$ and the stationary variance $\gamma^{2}=a\omega^{2}/2\beta$, which will be the parameter that we estimate. The exact transition density, i.e., the distribution of $\xi_{x+\Delta}$ given $\xi_{x}$ follows a non-central $\chi^{2}$ distribution scaled by a constant factor. Specifically, $\xi_{x+\Delta}|\xi_{x}\sim Z_{x}/2c$, where $Z_{x}\sim\chi^{2}_{\nu}(\lambda_{x})$ is non-central chi-squared distributed with $\nu=\frac{4a\beta}{\omega^{2}}$ degrees of freedom and non-centrality parameter $\lambda_{x+\Delta}=2c\rho\xi_{x}$. The multiplicative constant is $\frac{1}{2c}$, where $c=\frac{2\beta}{(1-\rho)\omega^{2}}$(Ditlevsen and Lansky, 2006).
@@
Our main goal is to infer $g(x,\theta)$ in order to date the observations $y_{i}$ to $t_{i}=g(x_{i},\theta),i=0,\ldots,n$. From $g(x,\theta)$ we are also able to predict the number of cycles in the signal:
Number of $\textrm{Number of Cycles}=\frac{g(x_{n},\theta)}{2\pi}.$ (5)
The signal (3) reaches peaks whenever $g(x,\theta)+b=\frac{\pi}{2}$ (modulo $\pi$). The amplitudes of these peaks can vary relative to the baseline value through $B(x)$. We will model these amplitude variations. To that end, assume that the signal has $K$ peaks and consider the set
$S=\left\{x_{k}^{\star}\in[0,n\Delta]\,\,\vline\,\,\exists n_{k}\in\mathbb{N}_{0}:g(x_{k}^{\star},\theta)+b=n_{k}\pi+\frac{\pi}{2}\right\}=\{x_{1}^{\star},\cdots,x_{K}^{\star}\}$ (6)
corresponding to all the high and low peaks in the compound signal (3) within the observation interval. The set is finite since $g(x,\theta)$ is strictly increasing with respect to $x$.
We model the amplitude process using a simple linear interpolation between signal peak values with locations in $S$. Let $B(x)=B_{0}+R(x)$ where $R(x)$ is a piecewise linear function with breaks at the peaks, $x^{*}_{1},\ldots,x^{*}_{K}$, and let $r_{1},\ldots,r_{K}$ be the respective differences between the peak amplitude and the baseline amplitude. Then we define $R(x)$ to be the linear interpolation of the $r_{k}$’s:
$R(x)=\begin{cases}r_{1}&x<x^{*}_{1}\\ \frac{r_{k+1}-r_{k}}{x^{*}_{k+1}-x^{*}_{k}}x+\frac{r_{k}x^{*}_{k+1}-r_{k+1}x^{*}_{k}}{x^{*}_{k+1}-x^{*}_{k}}&x^{*}_{k}\leq x\leq x^{*}_{k+1},k=1,\ldots,K-1\\ r_{K}&x>x^{*}_{K}\\ \end{cases}$ (7)
Note that $R(x_{k}^{*})=r_{k}$ so that $f(x_{k}^{*})=B_{0}+r_{k}\pm A$ in (3) where the sign is negative when $n_{k}$ is odd in $g(x_{k}^{*},\theta)+b=n_{k}\pi+\frac{\pi}{2}$ and positive when $n_{k}$ is even. Thus, $r_{k}$ captures any shift in amplitude relative to the baseline amplitude $B_{0}$, occurring at peak $k$. Since different signals exhibit varying number of peaks, the set of parameters $(r_{k})_{k\in 1,\ldots,K}$ has a variable length $K$, a quantity which is also estimated (see Section 3). The parameters of the model are summarized in Table 1.
@@
In Figure 1 we illustrate a signal simulated from the model (1) and (3), along with the embedded growth and amplitude processes (4) and (7).
@@
Figure 1: Example of a simulated signal. The dataset was simulated using $n=500$, $\Delta=1$, $\rho=0.82$, $\beta=0.2$, $\sigma=0.2$, $A=0.5$, $B_{0}=0.8$, $a=0.05$, $b=\frac{\pi}{2}$, $\omega^{2}=0.01$. The amplitude displacement parameters $r_{k}$ were simulated from a zero-mean normal distribution with variance 0.25. A. Simulated signal with measurement noise $y(x)$. B. Simulated hidden process $(\xi_{x})_{x\geq 0}$. C. Simulated growth-time process $g(x)$. D. Amplitude correction (relative to $B_{0}$) function $R(x)$. Gray dashed lines are signal peaks.
@@
$A$ $(0,\infty)$ Amplitude for the slow wave.
$B_{0}$ $(0,\infty)$ Baseline amplitude for the fast wave.
$b$ $(0,2\pi)$ Phase-offset of the signal.
$r_{k}$ $(A-B_{0},\infty)$ Displacements of the amplitudes from the baseline levels at peaks $k=1,\ldots,K$
@@
$a$ $(0,\infty)$ Mean of instantaneous growth process
$\beta$ $(0,\infty)$ Adjustment rate of instantaneous growth process
$\rho$ $(0,1)$ One-lag autocorrelation of instantaneous growth process, $\rho=\exp(-\Delta\beta)$
$\omega^{2}$ $(0,2a\beta)$ Infinitesimal variance of instantaneous growth process
$\gamma^{2}$ $(0,a^{2})$ Stationary variance of instantaneous growth process, $\gamma^{2}=a\omega^{2}/2\beta$

### v2

@@ Abstract
Signals with varying periodicity frequently appear in real-world phenomena, necessitating the development of efficient modelling techniques to map the measured nonlinear timeline to linear time. Here we propose a regression model that allows for a representation of periodic and dynamic patterns observed in time series data. The model incorporates a hidden strictly positive stochastic process that represents the instantaneous frequency, allowing the model to adapt and accurately capture varying time scales. A case study focusing on age estimation of narwhal tusks is presented, where cyclic element signals associated with annual growth layer groups are analyzed. We apply the methodology to data from one such tusk collected in West Greenland and use the fitted model to estimate the age of the narwhal. The proposed method is validated using simulated signals with known cycle counts and practical considerations and modelling challenges are discussed in detail. This research contributes to the field of time series analysis, providing a tool and valuable insights for understanding and modeling complex cyclic patterns in diverse domains.
@@
stochastic growth process,
@@ 1 Introduction
In some applications, the measurements are taken along a segment, where each measured point at some distance corresponds to an unknown time. Key examples are tree rings, sea sediment cores, ice cores, hair, ear stones (Monserud and Marshall, 2001; Jouzel and Masson-Delmotte, 2010; Charles et al., 1996; Hüssy et al., 2021a; Mosbacher et al., 2016), or in this case, narwhal tusks. A given variable is measured along a transect line at equidistant spatial points, and the oscillation patterns in the data are hypothesized to reflect yearly variations. The measurements can then be dated, i.e., distances can be transformed to a timeline. However, manual counting of cycles is often highly uncertain, and typically varies both between and within experts.
@@
A significant challenge presented by such datasets relates to the unknown nonlinear relation between distance and the underlying timeline, complicating the precise mapping of the signals to their actual temporal positions. In cases where reference signals are available or when the primary objective of the study centers on signal alignment rather than precise temporal mapping, methodologies like Graphical and Dynamic Time Warping (Jeong, Jeong and Omitaomu, 2011; Salvador and Chan, 2007; Wang et al., 2016) offer effective solutions for tasks involving pattern matching and classification. However, when reference signals are absent, and the exact mapping from spatial position to real time is of particular concern, the literature on time warping frameworks targeting cyclic signals is limited. One example is Dahlhaus et al. (2017), who developed a particle-EM variant that recovers the hidden phase and cycle shape jointly. Another example is Napolitano and Gardner (2016), who proposes two ways to estimate the time warping function in signals with irregular cycles, either directly estimating the warping functions or by using a probabilistic approach to construct a suitable estimator.
@@
This paper presents a novel approach for modelling noisy non-stationary cyclic signals with changing periodicity and amplitude. As opposed to the traditional seasonal decomposition (Dagum, 2013; Cleveland et al., 1990) where a signal is split into trend, seasonal, cyclic, and noise components, our model incorporates the seasonal and cyclic effects, and combine them into a single component. A key innovation in our model is the incorporation of a hidden strictly positive stochastic process, specifically the square-root diffusion process (Sørensen, 2012; Ditlevsen and Lansky, 2006), also known as the Cox-Ingersoll-Ross (CIR) process in the econometrics literature (Cox et al., 1985). The integral of this process is strictly increasing and models the material growth, by modulating the frequency while also respecting the direction of time. This provides realism in capturing the varying periodicity observed in real-world data.
@@
Estimation within our model is achieved using a stochastic variant of the EM algorithm, where we repeatedly filter the unobserved square-root diffusion process using Sequential Monte Carlo (SMC) and improve our estimates of the parameters in each new run. This procedure is explained in detail in section 3. While our model has applications across various domains, we focus here on its use in an ecological context. In this field, time series data are commonly associated with annual cycles, such as the growth patterns of organisms where various exogenous and endogenous factors can influence the speed of growth (Stounberg et al., 2022; Amais et al., 2021; Wooller et al., 2021). Our method can effectively capture the changing frequencies and provide more accurate predictions for these cyclic patterns.
We illustrate the method with data obtained from a narwhal (Monodon monoceros) tusk. The narwhal is a medium-sized odontocete (toothed whale), endemic to the Arctic regions. Understanding the age structure of wildlife populations is critical for assessments and conservation efforts (Watt et al., 2020). Age estimates form the basis of age-structured population models, which are essential tools for predicting demographic trends and informing management strategies, particularly for harvested or vulnerable species (Garde et al., 2022). Despite its importance, reliable age estimation in narwhals remains a challenge, and no fully accepted standard method has yet been established (Read, Hohn and Lockyer, 2018).
Narwhals are best known for the elongated, spiraled tusk of the male. Female narwhals rarely develop tusks; instead, they possess two small, embedded maxillary teeth. In males, the left maxillary tooth begins to erupt around 1–2 years of age and gradually elongates into the long tusk, which continues to grow throughout life. In contrast, the embedded teeth in both sexes generally cease growing between the ages of 10 and 20 years (Zhao et al., 2021; Watt et al., 2020).
Counting annual growth layer groups (GLGs)—each consisting of an opaque and a translucent layer—in dental tissues has traditionally been used to estimate the age of mammals (Hay, 1980; Read, Hohn and Lockyer, 2018). However, this method poses several challenges in narwhals (Hay, 1980; Zhao et al., 2021). For the embedded tooth, dentinal occlusion at sexual maturation obscures GLGs resulting in a minimum age estimate. For the erupted tusk, natural wear can gradually erode the tip of the tusk, removing early GLGs. This is mostly observed in older individuals, but will, as with the embedded tooth, result in a minimum age estimate. In addition, tusks seldom erupt in female narwhals and their large size, and the logistical and financial costs of collecting tusks limit its applicability in large-scale studies (Garde et al., 2012).
The most significant challenge, however, lies in the nature of the GLGs themselves: they are often faint, compressed or with subtle transitions between layers, making manual counting difficult. This process requires an experienced specialist, is inherently subjective and non-reproducible, and is associated with high variability both within and between observers.
Researchers have therefore explored two alternative techniques for age estimation. One promising approach is aspartic acid racemization (AAR), which has shown potential for estimating age in narwhals and other marine mammals (Garde et al., 2007, 2015; Watt et al., 2020). However, the initial development and calibration of the AAR method in narwhals relied on manual GLG counts from tusks (Garde et al., 2012), introducing uncertainty due to the inherent subjectivity of GLG interpretation. Additionally, AAR analysis requires access to the eye lens at the time of death, which can be logistically challenging to obtain. A recent study Garde et al. (2024) used radiocarbon (14C) dating to assess the accuracy of manual GLG-based age estimates in three tusks previously analyzed in Garde et al. (2012). This method leverages the so-called ”bomb pulse” - a sharp increase in atmospheric radiocarbon resulting from nuclear weapons testing in the mid- to late 1950s—which rapidly spread through global food webs, including marine ecosystems in the North Atlantic and Arctic Oceans. The bomb pulse serves as a temporal marker, allowing researchers to distinguish between pre- and post-pulse deposition of radiocarbon in narwhal tusks. However, this method is only of value for setting a benchmark for the bomb pulse and it can only be applied to whales born before or around the time of the bomb pulse. Radiocarbon dating results have shown that the manual GLG counts may overestimate the possible age. This is likely due to the challenges of reading tusks from older animals, for the reasons mentioned earlier. Our proposed model has the potential to reduce the subjectivity associated with manual GLG counting and to enhance our understanding of tusk growth in narwhals, such as the element composition in narwhal tusks , which may reflect underlying biological or environmental changes. In addition to its application in age estimation of narwhals and other toothed marine mammals, this method holds potential for broader use in any biological material that exhibits cyclic growth patterns.
In Section 2, we outline the model, in Section 3 we propose how to estimate model parameters and in Section 4 we perform a simulation study. In Section 5 we demonstrate the model and estimation technique in a narwhal tusk, where we focus on the mineralized elements (specifically the isotope Barium-137), which are deposited in the same annual pattern as the growth layers. For these data, decoding the hidden stochastic process of the model corresponds to uncovering the process that drives the growth of the tusk. Unlocking the growth process is of scientific interest, because it reveals life chapters of accelerating or declining growth, which may reflect underlying biological or environmental changes (Dietz et al., 2021). From this process, we are then able to estimate the number of cycles, thereby also the age of the narwhal. We conclude with a discussion of the model in section 6.
@@ 2 Time warping model
Consider a regression model with additive noise,
$y_{i}=f(x_{i},\theta)+\epsilon_{i},\quad i=0,1,\ldots,n,\,x_{i}=i\Delta,$ (1)
where $y_{i}$ is the measured variable (e.g., a chemical element) at spatial location $x_{i}$ and $\Delta=x_{i+1}-x_{i}$ is the distance between measurements. The observation interval is thus $0,n\Delta]$. The function $f$ is detailed below and depends on parameters $\theta$. The error terms $\epsilon_{i}$ are assumed independent and normally distributed with mean 0 and variance $\sigma^{2}$. The variable $x$ is the distance along a cross section of some object of interest, where a specific position $x_{i}$ corresponds to an unknown time through a monotonic and possibly non-linear function $g$,
$t_{i}=g(x_{i}).$ (2)
We impose that $g(x)$ is a strictly increasing function with respect to $x$, to ensure that time does not go backward.
For the signal, we assume linear combinations of sines and cosines to obtain oscillatory behavior. In particular, for the narwhal tusks, the data show clear indications of a large amplitude slow oscillation intertwined with lower amplitude oscillations of double frequency, with skewed peaks towards higher values. We therefore assume the form
$f(x;\theta)=A\sin(g(x)+b)-B\cos(2g(x)+2b),\quad g(0)=0,$ (3)
where $A>0$ and $B>0$ determine the amplitudes and $b$ is the phase at $x=0$. We require $A>B$ for identifiability, and also that $A/B$ is not too small - see supplementary material for a detailed discussion of both criteria. The signal (3) is composed of a low-frequency sine wave and a fast-frequency cosine wave: it traverses two cycles for each cycle of the sine wave. The sine wave may, for instance, represent annual cycles, while the fast cosine wave captures the summer-winter variations. See Figure 1 for an illustration. A generalization to higher frequency terms, for example, a model with higher order harmonics, is discussed in supplementary material.
There is no trend included in model (1), and we assume that it is removed in a prepossessing step, if necessary. For the signals used in our case study, no apparent trend was detected and only the average was subtracted. Furthermore, model (3) assumes constant amplitudes. We address this in Section 3.1.
Figure 1: Signal composition. Slow sine wave (blue) and fast cosine wave (orange) as well as their composition $f(x;\theta)$ (black), using $A=0.6$, $B=0.4$ and $b=0$ in Eq. (3). The growth process is non-stochastic and linear, $g(x)=x$.
In the linear and deterministic case where $g(x)=ax$ the cycles occur regularly (Figure 1). This is unrealistic in most settings, where the length of one cycle will typically vary due to environmental factors, leading to quasi-periodic dynamics. To describe this, we allow the instantaneous growth rate to be stochastic. We denote the unknown stochastic transformation $g(x)$ as the growth process. Let $g(x)=\int_{0}^{x}\xi_{s}ds$ where $(\xi_{x})_{x\geq 0}$ is a positive stochastic process. For $(\xi_{x})_{x\geq 0}$, we propose the square root diffusion process, satisfying the stochastic differential equation (SDE)
@@
where $W_{x}$ is a standard Wiener process. The hidden process $(\xi_{x})_{x\geq 0}$ describes instantaneous frequency deviations from the baseline frequency $a/2\pi$. The parameter $\beta$ determines the rate of adjustment to the baseline and $\omega$ scales the noise level. This is an ergodic process with stationary distribution the Gamma distribution with shape parameter $2\beta a/\omega^{2}$, scale parameter $\omega^{2}/(2\beta)$, mean $a$ and variance $a\omega^{2}/2\beta$, provided that $\beta,a,\omega>0$, and $2\beta a\geq\omega^{2}$ (Ditlevsen, Rubio and Lansky, 2020). These parameter restrictions ensure that the stochastic process generated by Eq. (4) is strictly positive with probability one (except at $x=0$) with long-term mean $a$, leading to a strictly increasing process $g(x)$. Denote the one-lag autocorrelation of the growth process by $\rho=\exp(-\Delta\beta)$ and the stationary variance by $\gamma^{2}=a\omega^{2}/2\beta$. The transition density, i.e., the distribution of $\xi_{x+\Delta}$ given $\xi_{x}$, follows a non-central $\chi^{2}$ distribution scaled by a constant factor. Specifically, $\xi_{x+\Delta}|\xi_{x}\sim Z_{x}/2c$, where $Z_{x}\sim\chi^{2}_{\nu}(\lambda_{x})$ is non-central chi-squared distributed with $\nu=4a\beta/\omega^{2}$ degrees of freedom, non-centrality parameter $\lambda_{x+\Delta}=2c\rho\xi_{x}$ and $c=2\beta/(1-\rho)\omega^{2}$ (Ditlevsen and Lansky, 2006).
@@
Our main goal is to infer the growth process $g(x)$ in order to date the observations $y_{i}$ to $t_{i}=g(x_{i})$, $i=0,\ldots,n$. From $g(x)$ we obtain the number of cycles in the signal:
$\textrm{Number of Cycles}=\frac{g(x_{n})}{2\pi}.$ (5)
@@
The parameters of the model are summarized in Table 1. In Figure 2 we illustrate a signal simulated from model (1) and (3), along with the embedded growth process $g(x)$ and the instantaneous frequency process generated by Eq. (4).
@@
Figure 2: Example of a simulated signal. The dataset was simulated using $n=500$, $\Delta=1$, $\beta=0.07$, $\sigma=0.3$, $A=0.6$, $B=0.4$, $a=0.05$, $b=\pi/20$, $\omega^{2}=0.064$. A. Simulated signal with measurement noise $y(x)$. B. Simulated hidden process $(\xi_{x})_{x\geq 0}$. C. Simulated growth process $g(x)$.
@@
AA $(0,\infty)$ Amplitude for the sine wave.
$B$ $(0,\infty)$ Amplitude for the cosine wave.
$b$ $0,2\pi)$ Phase-offset of the signal.
@@
$a$ $(0,\infty)$ Mean of instantaneous growth rate
$\beta$ $(0,\infty)$ Adjustment rate of instantaneous growth rate
$\rho$ $(0,1)$ One-lag autocorrelation of instantaneous growth rate, $\rho=\exp(-\Delta\beta)$
$\omega^{2}$ $(0,2a\beta)$ Infinitesimal variance of instantaneous growth rate
$\gamma^{2}$ $(0,a^{2})$ Stationary variance of instantaneous growth rate, $\gamma^{2}=a\omega^{2}/2\beta$